# Degree Distribution based Spiking Graph Networks for Domain Adaptation

### v3

@@ 4.4 Learning Framework
Finally, the overall training objective of DeSGDA integrates classification loss $\mathcal{L}_{S}$, adversarial training loss $\mathcal{L}_{AD}$, and pseudo-label distillation loss $\mathcal{L}_{T}$, which is formulated as:
@@
where $\lambda$ is a hyper-parameter to balance the adversarial training loss and classification loss. The learning procedure is illustrated in Algorithm D, and the complexity is shown in Appendix E .
@@ 5.1 Experimental Settings
Dataset. To demonstrate the effectiveness of DeSGDA, we conduct extensive experiments on four widely-used graph classification datasets from TUDataset  [^1] https://chrsmrrs.github.io/datasets/, including PROTEINS Dobson and Doig (2003), NCI1 Wale et al. (2008), FRANKENSTEIN Orsini et al. (2015), and MUTAGENICITY Kazius et al. (2005). To better address the variation in domain distributions within each dataset, we divided them into source and target domains based on the edge density, node density, and graph flux (i.e., the ratio of the number of nodes to the number of edges). The specific statistics, distribution visualization, and details introduction of experimental datasets are presented in Appendix F.
Baselines. We compare the proposed DeSGDA with competitive baselines on the aforementioned datasets, including one graph kernel method: WL subtree Shervashidze et al. (2011); four general graph neural networks: GCN Kipf and Welling (2017), GIN Xu et al. (2018), CIN Bodnar et al. (2021) and GMT Baek et al. (2021); two spiking graph neural networks: SpikeGCN Zhu et al. (2022) and DRSGNN Zhao et al. (2024); three domain adaptation methods: CDAN Long et al. (2018), ToAlign Wei et al. (2021b), and MetaAlign Wei et al. (2021a); and six graph domain adaptation methods: DEAL Yin et al. (2022b), CoCo Yin et al. (2023a), SGDA Qiao et al. (2023), DGDA Cai et al. (2024), A2GNN Liu et al. (2024a) and PA-BOTH Liu et al. (2024b). More detailed settings about the compared baselines are introduced in Appendix G.
@@
WL subtree 73.4 72.7 70.5 73.0 72.8 59.0 66.5 71.6 60.6 58.3 76.3 64.0 68.2
GCN 57.2$\pm$2.7 62.8$\pm$1.7 67.6$\pm$0.5 58.5$\pm$1.3 67.7$\pm$0.4 61.0$\pm$0.3 65.0$\pm$0.8 51.1$\pm$1.3 65.6$\pm$2.2 55.4$\pm$0.4 68.5$\pm$3.1 67.7$\pm$0.5 62.3
GIN 69.3$\pm$2.3 65.8$\pm$0.8 69.3$\pm$1.7 69.8$\pm$1.6 71.4$\pm$2.1 52.4$\pm$1.8 64.0$\pm$2.4 65.7$\pm$3.2 53.4$\pm$3.7 58.1$\pm$0.8 72.6$\pm$0.3 64.6$\pm$2.3 64.7
GMT 67.8$\pm$1.3 69.6$\pm$0.7 74.5$\pm$0.5 67.6$\pm$2.5 69.9$\pm$2.1 55.8$\pm$0.7 74.8$\pm$1.4 60.1$\pm$2.4 71.4$\pm$3.3 51.5$\pm$0.5 69.0$\pm$0.5 63.3$\pm$1.3 66.3
CIN 62.6$\pm$0.5 59.4$\pm$0.5 64.0$\pm$0.9 58.5$\pm$1.8 71.9$\pm$1.7 60.6$\pm$2.1 63.7$\pm$0.5 61.2$\pm$2.1 73.2$\pm$0.5 57.7$\pm$3.0 68.1$\pm$0.4 58.5$\pm$2.7 63.3
SpikeGCN 66.1$\pm$0.4 67.8$\pm$0.7 68.5$\pm$0.9 67.2$\pm$0.3 70.1$\pm$1.1 68.6$\pm$1.3 66.1$\pm$0.7 65.8$\pm$1.2 74.5$\pm$0.9 65.4$\pm$0.6 73.8$\pm$1.1 70.1$\pm$1.6 68.6
DRSGNN 68.3$\pm$0.5 71.9$\pm$0.9 71.5$\pm$1.1 75.1$\pm$1.7 76.7$\pm$1.3 74.4$\pm$0.8 70.5$\pm$0.7 67.8$\pm$0.8 77.0$\pm$1.5 65.9$\pm$1.6 75.2$\pm$1.9 69.2$\pm$2.1 71.8
CDAN 75.6$\pm$0.5 70.5$\pm$0.6 71.6$\pm$0.5 69.8$\pm$0.5 76.6$\pm$0.8 71.4$\pm$0.3 71.4$\pm$0.3 72.1$\pm$0.3 75.5$\pm$0.7 74.3$\pm$0.8 78.2$\pm$1.1 74.0$\pm$0.8 73.4
ToAlign 51.1$\pm$0.6 55.8$\pm$0.1 63.3$\pm$0.2 55.8$\pm$0.4 68.1$\pm$0.7 55.8$\pm$0.3 63.3$\pm$0.5 51.1$\pm$0.2 68.1$\pm$1.0 51.1$\pm$0.4 68.1$\pm$0.6 63.3$\pm$0.2 59.6
MetaAlign 59.4$\pm$1.1 62.2$\pm$1.0 68.9$\pm$0.3 65.3$\pm$0.8 75.1$\pm$0.7 67.5$\pm$2.1 70.9$\pm$1.4 60.6$\pm$2.3 72.4$\pm$1.4 59.4$\pm$0.6 74.6$\pm$0.7 67.8$\pm$1.3 67.0
DEAL 76.6$\pm$0.4 62.8$\pm$0.8 72.8$\pm$1.3 67.3$\pm$2.2 77.2$\pm$2.3 67.6$\pm$1.9 71.2$\pm$1.6 56.0$\pm$2.5 73.9$\pm$2.1 66.0$\pm$0.3 76.4$\pm$1.1 65.5$\pm$2.1 69.4
CoCo 73.4$\pm$0.5 73.6$\pm$0.8 73.4$\pm$1.0 71.6$\pm$0.5 75.2$\pm$1.6 74.6$\pm$0.3 70.7$\pm$0.8 68.4$\pm$1.5 75.0$\pm$0.2 72.7$\pm$0.4 76.3$\pm$1.1 75.0$\pm$1.8 73.3
SGDA 63.8$\pm$0.8 65.2$\pm$0.5 66.7$\pm$0.3 59.1$\pm$1.2 62.3$\pm$0.7 60.6$\pm$0.4 65.2$\pm$0.9 61.8$\pm$1.0 64.5$\pm$1.3 60.9$\pm$0.8 59.4$\pm$1.2 64.9$\pm$1.1 62.9
DGDA 59.4$\pm$0.7 62.3$\pm$1.1 63.1$\pm$0.5 61.2$\pm$0.9 60.4$\pm$0.6 58.8$\pm$1.0 60.3$\pm$0.8 63.5$\pm$1.2 61.9$\pm$0.8 60.4$\pm$1.6 64.2$\pm$1.3 62.6$\pm$1.4 61.5
A2GNN 65.4$\pm$0.7 66.4$\pm$1.1 65.7$\pm$1.3 66.0$\pm$0.6 64.9$\pm$1.2 65.8$\pm$1.6 65.5$\pm$1.8 66.0$\pm$1.4 65.8$\pm$2.1 65.6$\pm$1.9 66.1$\pm$1.7 66.0$\pm$2.0 65.8
PA-BOTH 66.9$\pm$0.5 67.1$\pm$0.8 67.3$\pm$1.1 65.8$\pm$0.7 69.1$\pm$1.0 66.1$\pm$1.4 66.7$\pm$1.3 67.4$\pm$1.4 66.3$\pm$1.8 66.0$\pm$1.2 66.8$\pm$0.8 66.3$\pm$1.5 66.8
DeSGDA 78.7$\pm$1.3 78.4$\pm$1.1 74.8$\pm$0.6 77.6$\pm$0.9 79.5$\pm$1.2 76.7$\pm$0.8 74.9$\pm$0.7 71.2$\pm$1.7 79.5$\pm$1.4 72.8$\pm$0.8 81.0$\pm$1.5 75.1$\pm$1.0 76.7
Implementation Details. DeSGDA and all baseline models are implemented using PyTorch [^2] https://pytorch.org/ and PyTorch Geometric [^3] https://www.pyg.org/. For DeSGDA, we deploy the GIN Xu et al. (2018) as the backbone of the degree-aware personalized spiking graph encoder, incorporating a mean-pooling layer for the readout function. We conduct experiments for DeSGDA and all baselines on NVIDIA A100 GPUs for a fair comparison, where the learning rate of Adam optimizer set to ${10^{-4}}$, hidden embedding dimension 256, weight decay ${10^{-12}}$, and GNN layers 4. Additionally, DeSGDA and all baseline models are trained using all labeled source samples and evaluated on unlabeled target samples Wu et al. (2020). The performances of all models are measured and averaged on all samples for five different runs.
Table 2: The results of ablation studies on PROTEINS (source → target). Bold results indicate the best performance.
[…]
DeSGDA w/o CA 72.3 72.6 71.9 75.1 74.1 71.3 71.5 70.4 71.6 70.5 78.3 71.5 72.6
DeSGDA w/o PL 72.1 68.7 67.8 69.4 63.7 55.7 68.3 69.5 70.2 69.6 76.7 66.9 68.3
DeSGDA w/o CF 56.7 54.2 60.7 62.1 73.4 68.3 55.3 69.5 76.7 65.5 73.7 63.3 65.0
DeSGDA w/o TL 61.7 43.8 56.1 71.2 49.5 69.0 42.4 68.3 73.7 64.0 70.5 43.5 59.5
DeSGDA w/ PT 71.5 68.3 66.1 72.4 66.7 70.1 67.9 70.8 66.3 71.8 74.1 72.9 70.0
DeSGDA w/ CL 74.9 73.3 73.7 75.1 77.0 71.3 73.9 70.0 78.1 77.6 78.8 74.9 74.9
DeSGDA 78.7 78.4 74.8 77.6 79.5 76.7 74.9 71.2 79.5 72.8 81.0 75.1 76.7
5.2 Performance Comparision
We present the results of our proposed DeSGDA with all baseline models under the setting of graph domain adaptation on different datasets in Tables 1 and 13-23. From these tables, we observe that: (1) The performance of graph domain adaptation methods surpasses graph-based and spike-based graph methods. We attribute this improvement to the fact that domain distribution shifts degrade the performance of traditional graph methods. (2) The graph domain adaptation methods (DEAL and CoCo) outperform the spike-based graph methods (SpikeGCN and DRSGNN), underscoring the necessity of the research in spiking graph domain adaptation. (3) The WL subtree method outperforms SGDA, DGDA, A2GNN, and PA-BOTH but falls short compared to DEAL and CoCO. We attribute this to the relatively limited research specifically addressing the graph domain adaptation problem (e.g., DEAL and CoCo). To bridge this gap, we adapted node classification methods for graph classification tasks (e.g., SGDA, DGDA, A2GNN, and PA-BOTH). While the WL subtree method demonstrates superior performance over these adapted node classification-based methods, it remains less effective than dedicated graph domain adaptation methods tailored for graph classification tasks. (4) Our DeSGDA outperforms all baselines for most cases, which demonstrates its superiority over other methods. The remarkable performance of DeSGDA lies in two main reasons: (i) The degree-aware personalized spiking representations can capture more expressive information for graph classification by dynamically adjusting the thresholds of nodes in SNNs. (ii) The adversarial distribution alignment effectively addresses domain discrepancies by adversarially training the encoder and domain discriminator to align feature spaces. Moreover, the pseudo-label distillation aids in updating the source degree thresholds in the target domain, thereby ensuring optimal performance. More results on other datasets can be found in Appendix H.1.
Figure 2: The performance with different GNN architectures on PROTEINS.
Additionally, we conduct experiments to explore the flexibility of DeSGDA. Specifically, we replace the backbone of the degree-aware personalized spiking graph encoder (GIN) with various GNN methods (i.e., GCN and GAT), and the results are shown in Figure 2. From those results, we observe that GIN consistently outperforms other GNN architectures in most cases, demonstrating its powerful representation capability. This phenomenon also justifies our choice of using GIN to enhance the performance of the proposed DeSGDA. More results are reported in Appendix H.1.

### v4

@@ 4.4 Learning Framework
Overall, the training objective of DeSGraDA integrates classification loss $\mathcal{L}_{S}$, temporal-based distribution alignment loss $\mathcal{L}_{AD}$, and pseudo-label distillation loss $\mathcal{L}_{T}$, which is formulated as:
@@
where $\lambda$ is a hyper-parameter to balance the distribution alignment loss and classification loss. The learning procedure is illustrated in Algorithm F, and the complexity is shown in Appendix G.
[…]
Methods SEED BCI
[…]
WL subtree 40.7 36.8 42.6 35.0 35.6 38.3 47.9 47.7 46.0 46.7 47.7 47.5
GCN 46.5$\pm$0.6 47.4$\pm$0.9 46.6$\pm$1.2 47.7$\pm$1.4 45.8$\pm$1.2 47.1$\pm$1.6 49.6$\pm$2.5 48.7$\pm$2.8 51.1$\pm$1.0 51.5$\pm$2.0 49.6$\pm$2.4 49.1$\pm$1.7
GIN 47.4$\pm$1.7 48.0$\pm$1.6 47.3$\pm$1.4 47.5$\pm$1.8 41.6$\pm$2.0 46.1$\pm$1.3 49.4$\pm$2.5 48.4$\pm$2.1 51.8$\pm$1.4 51.2$\pm$2.5 50.0$\pm$1.7 48.7$\pm$2.1
GMT 46.5$\pm$0.5 47.8$\pm$1.3 47.2$\pm$0.7 47.2$\pm$1.3 46.4$\pm$0.9 46.4$\pm$1.2 48.8$\pm$1.3 47.8$\pm$1.1 49.4$\pm$1.0 48.5$\pm$1.5 50.7$\pm$0.9 51.5$\pm$1.5
CIN 46.9$\pm$0.5 48.4$\pm$1.1 47.0$\pm$1.5 47.3$\pm$0.7 47.0$\pm$1.6 47.0$\pm$0.9 50.3$\pm$1.6 48.8$\pm$1.5 50.4$\pm$1.3 50.4$\pm$1.2 50.1$\pm$1.5 50.9$\pm$1.7
SpikeGCN 46.3$\pm$1.0 47.4$\pm$0.8 45.8$\pm$1.2 47.7$\pm$1.1 45.8$\pm$1.5 46.4$\pm$1.2 52.5$\pm$1.6 52.8$\pm$1.3 54.1$\pm$1.9 52.1$\pm$1.0 51.7$\pm$1.8 50.5$\pm$2.3
DRSGNN 47.1$\pm$1.0 48.5$\pm$0.9 46.5$\pm$1.2 48.1$\pm$1.3 46.9$\pm$0.8 47.6$\pm$1.4 52.7$\pm$1.3 52.8$\pm$1.9 53.8$\pm$1.5 52.7$\pm$1.6 53.0$\pm$2.1 51.3$\pm$1.8
CDAN 52.6$\pm$1.2 54.5$\pm$0.7 53.9$\pm$0.7 55.9$\pm$1.3 51.6$\pm$1.1 53.6$\pm$0.8 51.9$\pm$1.3 52.6$\pm$1.4 51.8$\pm$1.1 55.4$\pm$1.8 52.5$\pm$1.5 53.1$\pm$1.4
ToAlign 51.2$\pm$1.3 52.3$\pm$0.8 51.5$\pm$0.9 49.7$\pm$1.5 49.6$\pm$1.1 49.4$\pm$1.3 52.5$\pm$1.7 53.7$\pm$1.5 52.2$\pm$1.4 54.4$\pm$1.2 52.7$\pm$1.0 51.8$\pm$1.3
MetaAlign 51.2$\pm$1.4 53.7$\pm$0.9 52.2$\pm$1.1 53.8$\pm$0.8 51.2$\pm$1.4 52.0$\pm$1.2 51.1$\pm$1.5 51.8$\pm$1.2 50.4$\pm$1.7 52.5$\pm$1.5 51.7$\pm$1.5 51.3$\pm$1.1
DEAL 57.4$\pm$1.1 57.5$\pm$1.4 56.6$\pm$0.7 58.1$\pm$1.2 53.9$\pm$0.7 57.8$\pm$1.3 53.7$\pm$1.4 52.5$\pm$2.2 52.6$\pm$1.6 54.5$\pm$1.4 52.7$\pm$1.7 52.8$\pm$1.2
CoCo 55.5$\pm$1.5 56.7$\pm$0.7 56.3$\pm$1.3 58.8$\pm$0.8 54.2$\pm$1.2 57.5$\pm$1.3 54.0$\pm$1.3 55.2$\pm$2.5 52.7$\pm$2.1 52.7$\pm$1.9 51.7$\pm$2.8 51.0$\pm$2.4
SGDA 47.1$\pm$0.6 41.6$\pm$1.4 43.8$\pm$0.7 45.9$\pm$1.2 49.4$\pm$1.1 50.1$\pm$1.5 49.7$\pm$1.6 48.4$\pm$1.5 50.6$\pm$1.0 50.4$\pm$1.3 50.5$\pm$1.2 50.7$\pm$1.4
StruRW 47.1$\pm$0.9 45.9$\pm$0.7 46.5$\pm$1.3 48.2$\pm$1.2 46.9$\pm$1.2 47.3$\pm$1.4 48.7$\pm$1.1 47.3$\pm$1.7 49.5$\pm$1.1 49.7$\pm$1.5 50.0$\pm$1.8 50.2$\pm$1.6
A2GNN 47.6$\pm$1.2 47.6$\pm$0.9 46.2$\pm$0.8 48.3$\pm$1.1 46.2$\pm$1.0 47.9$\pm$0.6 52.0$\pm$1.7 53.0$\pm$1.4 52.0$\pm$1.0 53.7$\pm$1.1 52.2$\pm$1.3 51.8$\pm$1.7
PA-BOTH 48.2$\pm$1.4 48.2$\pm$0.8 47.3$\pm$1.2 48.3$\pm$1.0 48.5$\pm$1.2 45.2$\pm$0.6 49.2$\pm$1.6 50.0$\pm$1.2 51.1$\pm$1.3 51.3$\pm$1.5 50.5$\pm$1.6 48.8$\pm$1.4
DeSGraDA 58.0$\pm$1.5 58.2$\pm$1.4 57.0$\pm$1.8 58.3$\pm$1.6 55.9$\pm$2.1 58.1$\pm$1.6 54.1$\pm$1.5 53.6$\pm$1.6 54.9$\pm$1.1 56.2$\pm$1.8 55.0$\pm$1.3 54.6$\pm$1.2
@@ 5.1 Experimental Settings
Dataset. To evaluate the effectiveness of DeSGraDA, we conduct extensive experiments across two types of domain shifts: (1) structure-based domain shifts, where the discrepancy between domains arises primarily from differences in graph topology, such as variations in node and edge densities. This category includes datasets DD, PROTEINS (Dobson and Doig, 2003), SEED Zheng and Lu (2015); Duan et al. (2013), and BCI (Brunner et al., 2008); (2) feature-based domain shifts, where domains differ mainly in semantic information while maintaining similar graph structures. This setting involves datasets including DD, PROTEINS, BZR, BZR_MD, COX2, and COX2_MD (Dobson and Doig, 2003; Sutherland et al., 2003). The specific statistics, distribution visualization, and detailed introduction of experimental datasets are presented in Appendix H.
Baselines. We compare the proposed DeSGraDA with competitive baselines on the aforementioned datasets, including one graph kernel method: WL subtree (Shervashidze et al., 2011); four graph-based methods: GCN (Kipf and Welling, 2017), GIN (Xu et al., 2018), CIN (Bodnar et al., 2021) and GMT (Baek et al., 2021); two spiking-based graph methods: SpikeGCN (Zhu et al., 2022) and DRSGNN (Zhao et al., 2024); three domain adaptation methods: CDAN (Long et al., 2018), ToAlign (Wei et al., 2021b), and MetaAlign (Wei et al., 2021a); and six graph domain adaptation methods: DEAL (Yin et al., 2022), CoCo (Yin et al., 2023), SGDA (Qiao et al., 2023), StruRW (Liu et al., 2023), A2GNN (Liu et al., 2024a) and PA-BOTH (Liu et al., 2024b). More settings about baselines are introduced in Appendix I and J.
Table 2: Graph classification results (in %) under node and edge density domain shifts on the PROTEINS dataset, and feature domain shifts on DD, PROTEINS, BZR, BZR_MD, COX2, and COX2_MD. For convenience, PROTEINS, DD, COX2, COX2_MD, BZR, and BZR_MD are abbreviated as P, D, C, CM, B, and BM, respectively. Bold results indicate the best performance.
Methods Node Shift Edge Shift Feature Shift
@@
WL subtree 69.1 61.2 41.6 68.7 50.7 58.1 43.0 42.2 53.1 58.2 51.3 44.0
GCN 73.7$\pm$0.3 57.6$\pm$0.2 24.4$\pm$0.4 73.4$\pm$0.2 57.6$\pm$0.2 24.0$\pm$0.1 48.9$\pm$2.0 60.9$\pm$2.3 51.2$\pm$1.8 66.9$\pm$1.8 48.7$\pm$2.0 78.8$\pm$1.7
GIN 71.8$\pm$2.7 58.5$\pm$4.3 74.2$\pm$1.7 62.5$\pm$4.7 53.0$\pm$4.6 73.7$\pm$0.8 57.3$\pm$2.2 61.9$\pm$1.9 53.8$\pm$2.5 55.6$\pm$2.0 49.9$\pm$2.4 79.2$\pm$2.8
GMT 73.7$\pm$0.2 57.6$\pm$0.3 75.6$\pm$1.4 73.4$\pm$0.3 57.6$\pm$0.1 24.0$\pm$0.1 59.5$\pm$2.5 50.7$\pm$2.2 49.3$\pm$1.8 58.2$\pm$2.0 50.2$\pm$2.3 74.4$\pm$1.8
CIN 74.1$\pm$0.6 60.1$\pm$2.1 75.6$\pm$0.2 74.5$\pm$0.2 57.8$\pm$0.2 75.6$\pm$0.6 59.1$\pm$2.6 58.0$\pm$2.7 51.2$\pm$2.0 55.6$\pm$1.5 49.2$\pm$1.4 74.2$\pm$1.9
SpikeGCN 71.8$\pm$0.9 64.9$\pm$1.4 71.1$\pm$1.9 71.8$\pm$0.8 63.8$\pm$1.0 68.6$\pm$1.1 59.6$\pm$2.2 63.3$\pm$1.8 52.6$\pm$2.5 68.6$\pm$1.8 53.3$\pm$1.7 76.1$\pm$2.0
DRSGNN 73.6$\pm$1.1 64.6$\pm$1.2 70.2$\pm$1.7 72.6$\pm$0.6 63.1$\pm$1.4 70.4$\pm$1.9 60.9$\pm$2.4 65.4$\pm$1.9 52.9$\pm$1.8 66.9$\pm$2.3 52.8$\pm$1.7 76.4$\pm$2.7
CDAN 75.9$\pm$1.0 60.8$\pm$0.6 75.8$\pm$0.3 72.2$\pm$1.8 59.8$\pm$2.1 69.3$\pm$4.1 59.4$\pm$2.0 63.1$\pm$2.7 51.2$\pm$2.3 68.2$\pm$1.8 50.7$\pm$1.6 75.2$\pm$1.9
ToAlign 73.7$\pm$0.4 57.6$\pm$0.6 24.4$\pm$0.1 73.4$\pm$0.1 57.6$\pm$0.1 24.0$\pm$0.3 62.1$\pm$2.1 66.5$\pm$2.3 53.2$\pm$2.6 55.8$\pm$2.3 56.2$\pm$3.0 78.8$\pm$2.4
MetaAlign 74.3$\pm$0.8 60.6$\pm$1.7 76.3$\pm$0.3 75.5$\pm$0.9 64.8$\pm$1.6 69.3$\pm$2.7 63.3$\pm$2.1 66.2$\pm$1.9 51.2$\pm$2.0 69.5$\pm$2.3 48.7$\pm$1.8 76.8$\pm$2.7
DEAL 75.4$\pm$1.2 68.1$\pm$1.9 73.8$\pm$1.4 76.5$\pm$0.4 67.5$\pm$1.3 76.0$\pm$0.2 70.6$\pm$1.9 66.8$\pm$2.5 50.9$\pm$2.4 67.8$\pm$1.9 51.1$\pm$2.3 79.4$\pm$2.2
CoCo 74.8$\pm$0.6 65.5$\pm$0.4 72.4$\pm$2.9 75.5$\pm$0.2 59.8$\pm$0.5 73.6$\pm$2.3 66.0$\pm$2.7 61.2$\pm$2.3 53.6$\pm$1.8 78.2$\pm$2.0 57.8$\pm$1.6 79.8$\pm$1.8
SGDA 64.2$\pm$0.5 66.9$\pm$1.2 65.4$\pm$1.6 63.8$\pm$0.6 66.7$\pm$1.0 60.1$\pm$0.8 48.3$\pm$2.0 55.8$\pm$2.6 49.8$\pm$1.8 66.9$\pm$2.3 50.3$\pm$2.1 78.8$\pm$2.6
A2GNN 65.7$\pm$0.6 66.3$\pm$0.9 65.2$\pm$1.4 65.4$\pm$1.3 68.2$\pm$1.4 65.4$\pm$0.7 57.8$\pm$2.1 60.3$\pm$1.5 51.5$\pm$1.8 67.7$\pm$2.1 51.6$\pm$2.3 77.5$\pm$1.9
StruRW 71.9$\pm$2.3 66.7$\pm$1.8 52.8$\pm$1.9 72.6$\pm$2.2 66.2$\pm$2.2 48.9$\pm$2.0 59.1$\pm$2.3 58.8$\pm$2.8 51.2$\pm$2.0 54.8$\pm$2.9 49.2$\pm$1.4 74.7$\pm$2.1
PA-BOTH 61.0$\pm$0.8 60.3$\pm$0.6 63.7$\pm$1.5 63.1$\pm$0.7 64.3$\pm$0.5 66.3$\pm$0.7 54.2$\pm$3.2 56.7$\pm$2.6 52.9$\pm$2.8 61.8$\pm$2.0 47.5$\pm$3.0 78.8$\pm$1.9
DeSGraDA 76.3$\pm$1.9 69.2$\pm$2.3 77.5$\pm$2.2 76.8$\pm$1.9 68.6$\pm$1.8 76.5$\pm$2.8 73.6$\pm$1.9 71.2$\pm$1.6 54.7$\pm$1.8 78.6$\pm$2.2 56.3$\pm$1.5 80.3$\pm$1.9
5.2 Performance Comparison
Figure 3: Correlation comparisons of spiking frequency and aggregation weights under adaptive thresholds on PROTEINS and DD datasets.
We present the results of the proposed DeSGraDA with all baselines under two types of domain shifts on different datasets in Tables 1, 2, and 11-14. From these tables, we observe that: (1) GDA methods outperform traditional graph-based and spiking-based graph methods in most cases, highlighting the adverse impact of domain distribution shifts on conventional approaches and underscoring the importance of advancing research in spiking graph domain adaptation. (2) The spiking-based graph methods (i.e., SpikeGCN and DRSGNN) outperform the models specific for node classification (i.e., SGDA, StruRW, A2GNN, and PA-BOTH) but fall short compared to models for graph classification (i.e., DEAL and CoCo). This performance gap is primarily due to the limited exploration of graph classification under domain shift. Although spiking-based methods exhibit advantages over adapted node classification models, they remain less effective than specialized graph domain adaptation methods explicitly designed for graph-level tasks. (3) DeSGraDA outperforms all baselines in most cases, demonstrating its advantage over other methods. The superior performance can be attributed to two key factors. First, the degree-conscious spiking representations dynamically adjust node-specific firing thresholds in SNNs, enabling the model to capture more expressive and discriminative graph features. Second, the temporal-based distribution alignment aligns source and target domain representations by matching spiking membrane dynamics over time, effectively mitigating distributional discrepancies. Moreover, the pseudo-label distillation helps refine degree thresholds in the target domain, further enhancing generalization. More results can be found in Appendix K.1.
We further conduct experiments to examine the correlation between spiking frequency and aggregation weights under adaptive thresholds across different node degrees. As shown in Figure 3, the correlation coefficient under adaptive thresholds is significantly lower than that under fixed thresholds, demonstrating that DeSGraDA effectively smooths spiking frequency across node degrees, mitigates over-activation in high-degree nodes, and promotes balanced information aggregation.
Table 3: The results of ablation studies on the PROTEINS dataset (source → target). Bold results indicate the best performance.
@@ 5.2 Performance Comparison
DeSGraDA w/ CDAN 73.3 82.4 67.8 76.5 74.4 78.7 66.5 71.3 73.7 70.2 74.1 68.8
DeSGraDA w/o PL 73.7 81.2 67.1 81.2 75.9 79.6 67.8 71.9 74.7 69.0 75.4 68.3
DeSGraDA w/o CF 65.0 67.4 53.5 64.5 66.6 68.8 56.7 63.4 66.1 60.9 53.9 56.6
DeSGraDA w/o TL 73.6 80.6 65.6 80.6 73.1 78.4 63.9 69.3 69.6 68.7 72.9 64.5
DeSGraDA 76.3 84.6 69.2 83.6 77.5 83.7 69.8 74.0 76.2 73.0 77.8 70.5
5.3 Ablation Study
We conduct ablation studies to examine the contributions of each component: (1) DeSGraDA w/ CDAN: It utilizes the static distribution alignment instead of temporal-based module; (2) DeSGraDA w/o PL: It removes the pseudo-label distilling module; (3) DeSGraDA w/o CF: It removes the classification loss $\mathcal{L}_{S}$; (4) DeSGraDA w/o TL: It utilizes the fixed global thresholds on all nodes.
Experimental results are shown in Table 3. From the table, we find that: (1) The degree-aware thresholding mechanism substantially improves representational capacity. When replaced with fixed global thresholds (DeSGraDA w/o TL), performance consistently declines, suggesting that dynamically adjusting thresholds based on node degree helps the model better handle structural heterogeneity and encode informative spiking representations. (2) The temporal-based distribution alignment and pseudo-label distillation modules are crucial for effective domain adaptation. Removing the temporal alignment module (DeSGraDA w/ CDAN) leads to consistent performance drops across most tasks, highlighting its role in mitigating domain shifts by aligning membrane dynamics across domains. Similarly, eliminating pseudo-label distillation (DeSGraDA w/o PL) degrades performance, demonstrating the importance of leveraging confident target predictions to refine threshold adaptation and support generalization. (3) DeSGraDA outperforms all ablated variants across all domain shifts, confirming the complementary strengths of its core components. Notably, removing the classification loss (DeSGraDA w/o CF) results in the largest performance degradation, highlighting the necessity of source supervision for learning discriminative features. Furthermore, we provide ablation studies to examine the effect of directly replacing SGNs with commonly used GNNs for DeSGraDA, and the results are shown in Tables 7, 8. More ablation studies are reported in Appendix K.3.